OpenStudy (anonymous):

log(5)16 – log(5)2t = log(5)2 Please help. I do not understand at all!!

OpenStudy (ankit042):

you know log A- logB =log(A/B)...now you can move forward

OpenStudy (anonymous):

So log(5)16 / log(5)2t = og(5)16 – log(5)2t??

OpenStudy (ankit042):

NO it will be log(5)(16/2t)= log(5)16 – log(5)2t

OpenStudy (anonymous):

Oh.

OpenStudy (ankit042):

Now can you complete from here?

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

Wait, once you set it all up, can the logs cancel out?

OpenStudy (ankit042):

yes if log are of same base you can remove them! Hope this helps

OpenStudy (anonymous):

Would it be log(5)16/2t = log(5)2 16/2t = 2 16 = 2*2t 8 = 2t 4=t??? @ankit042

OpenStudy (ankit042):

Yes you are right!

OpenStudy (anonymous):

Okay, thanks!